Solve 3^(2x+1)=2^(5x-2) to the nearest tenth
Answer by HyperBrain(694)

32x+1=25x-2
Step 1: Find the logaithms of both sides. Note that log(ab)=b log a
(2x+1)log 3=(5x-2)log 2
0.47712(2x+1)=0.30102(5x-2)
1.58501(2x+1)=5x-2
3.17002x+1.58501=5x-2
3.58501=1.82997x
x=1.95905=2.0 (rounded to the nearest tenth)
::
Power up,
HyperBrain!
::
http://adrianvidal.wordpress.com
http://anothervidal.wordpress.com